OpenStudy (mitu12):

solve for g in the equation: S=1/2gt^2 i am really confused and i dont know where to begin pleaseee help!!!!

OpenStudy (anonymous):

is it \[s=\frac{1}{2\cdot g \cdot t^2} ?\]

OpenStudy (wolf1728):

I believe that's the equation for the distance fallen of a dropped object. So it is distance = .5*g*t^2 The value of 'g' is 9.8 meters per sec^2

OpenStudy (anonymous):

or \[\frac{1}{2}\cdot g \cdot t^2\]

OpenStudy (wolf1728):

that's it bakur

OpenStudy (mitu12):